OpenStudy (anonymous):

(2 x^1/2y)/(xy^3)^3/2 I Don't Understand This Please Help

OpenStudy (anonymous):

do you need to find the variable or answer the equation?

OpenStudy (anonymous):

^no

OpenStudy (anonymous):

no to wat?

OpenStudy (anonymous):

you have to simplify it

OpenStudy (anonymous):

you cant find values of x and y from 1 equation

OpenStudy (anonymous):

Oh ok...

OpenStudy (anonymous):

Sorry not good with simplifying answers. :(

OpenStudy (anonymous):

can you write the problem using equation lana

OpenStudy (anonymous):

or maybe capture a pic n post

OpenStudy (anonymous):

@ganeshie8 y is in power x's power i dont see any good simplification for this

OpenStudy (anonymous):

or she wrote something wrong

ganeshie8 (ganeshie8):

lol the expression looks flattened out -_- maybe something like below : \(\large \dfrac{2 x^{\frac{1}{2}}y}{(xy^3)^{\frac{3}{2}}} \)

OpenStudy (anonymous):

\[\frac{ 2x^{1/2y} }{ xy^{3(3/2)} }\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

this makes sense :P

ganeshie8 (ganeshie8):

haha oly @LannaBoo24 can conform :)

OpenStudy (anonymous):

Sorry You Guys I Sorta Turned It In But Thanks For The Help But Yes It Was Simplifying.Also Id Love Some Help On My Homework If One Of You Could?